From the values of f(x) = xex given in the table
x 1.8 1.9 2.0 2.1 2.2
f(x) 10.8894 12.7032 14.7787 17.1489 19.8550
find f
00(2.0) using the central difference formula of O(h
2
) for h = 0.1 and h = 0.2.
Calculate T.E. and actual error.
1
Expert's answer
2020-04-23T18:24:23-0400

x 1.8 1.9 2.0 2.1 2.2

f(x) 10.8894 12.7032 14.7787 17.1489 19.8550

to find f''(2.0) using Central Difference Formula


"f'(x)=[f(x+h)-f(x-h)]\/2h" or"f'(x)=[f(x+h)-f(x)]\/h"

"f(x)=xe^2\\

f'(x)=xe^x+e^x\\
f''(x)=2e^x+xe^x\\
f'''(x)=3e^x+xe^x\\
f^4(x)=4e^x+xe^x"

given "x=2.0 ;h=0.1; h=0.2"

we have

"f''(x)=[f(x+h)-2f(x)+f(x-h)]\/h^2"

hence

"f''(2.0)=[f(2.1)-2f(2.0)+f(1.9)]/0.1^2\\
f''(2.0)=[17.1489-2(14.7787)+12.7032]/0.01\\
f''(2.0)=29.47"


finding the approximated error (T.E)

error at h=0.1

"O(h^2)=(h^2/12)|-f^4(h)|\\
=(0.01/12)|-(4e^x+xe^x)|\\
=(0.01/12)|-(4.4206837+0.1105171)|\\
=0.003776"


error at h=0.01

"O(h^2)=(h^2/12)|-f^4(h)|\\
=(0.0001/12)|-(4.0402007+0.0101005)|\\
=(0.0001/12)|-4.0503012|\\
=0.000033752"

to find actual error

at h=0.1

"=|1-[he^h+h*exp(-h)]/h^2|\\
=|1-[0.1105171+0.0904837]/0.01|\\
=19.10008"

at h=0.01

"=|1-[he^h+h*exp(-h)]/h^2|\\
=|1-[0.0101005+0.009900498]/0.0001|\\
=199.00998"